OpenStudy (anonymous):

A fair die is tossed 4 times. Find the probability of a. at least 2 5s b. at most 3 even numbers c. no more than 1 odd number d. at least 1 prime number e. at most 1 number greater than 4

OpenStudy (yuki):

you use the fact that P(A) = 1- P(not A)

OpenStudy (anonymous):

can u show me how to do each part?

OpenStudy (yuki):

let's do a

OpenStudy (yuki):

P(at least two 5s) = 1- P(not at least two 5s)

OpenStudy (yuki):

= 1- P(at most one 5)

OpenStudy (yuki):

so let's find P(at most one 5)

OpenStudy (yuki):

can you find this ?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

I got 146/1296 for A

OpenStudy (anonymous):

is that right???

OpenStudy (yuki):

I am not finding the answer, just trying to help

OpenStudy (yuki):

so P(at most one 5) = P(no 5s) + P(exactly one 5)

OpenStudy (yuki):

P(no 5) = \[(5/6)^4\]

OpenStudy (yuki):

P(exactly one 5) follows the binomial distribution, so \[4C1*(1/5)^1*(5/6)^4\]

OpenStudy (yuki):

when you subtract the sum of those from 1, you get the answer for a),

OpenStudy (anonymous):

wait

OpenStudy (yuki):

oops, the 1/5 should be 1/6

OpenStudy (anonymous):

don't u need 0 5s as well?

OpenStudy (yuki):

I already did that, it is (5/6)^4

OpenStudy (anonymous):

wait question is at least 25s

OpenStudy (yuki):

let me explain you that par again

OpenStudy (yuki):

P( at least two 5s) = P(exactly two) + P(exactly three)+P(exactly four)

OpenStudy (anonymous):

oh yea I did that

OpenStudy (yuki):

it is a pain to calculate all those, so we use the rule P(event) = 1-P(complement of that event)

OpenStudy (anonymous):

4C2(1/6)^2(5/6)^3 + 4C3(1/6)^3(5/6)^1 + 4C4(1/6)^4(5/6)^0

OpenStudy (anonymous):

is that right?

OpenStudy (yuki):

Yes.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

B

OpenStudy (yuki):

what I did was 1- (5/6)^6 -4C1(1/6)(5/6)^3

OpenStudy (yuki):

for B, the complement is "no evens at all"

OpenStudy (yuki):

so 1-P(no evens) will be the answer

OpenStudy (anonymous):

isn't it \[4C3(1/2)^3(1/2)^1 + 4C2(1/2)^2(1/2)^2 + 4C1 (1/2)^1(1/2)^3\]

OpenStudy (anonymous):

???

OpenStudy (yuki):

you need to include 4C0(1/2)^4

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

everything else is correct?

OpenStudy (yuki):

az, just to let you know you probably know how to get the answers for all of this. But you also need to know that you are missing the point of what these problems are asking.

OpenStudy (yuki):

Yes.

OpenStudy (anonymous):

next one idk

OpenStudy (anonymous):

so it can be 0 or 1 odd number?

OpenStudy (anonymous):

??

OpenStudy (yuki):

yep

OpenStudy (anonymous):

4C1(1/2)^1 (1/2)^3 + 4C0(1/2)^0 (1/2)^4 ??

OpenStudy (anonymous):

is that right

OpenStudy (yuki):

yes

OpenStudy (yuki):

az, I need to go soon so let me leave you an advice

OpenStudy (yuki):

your method is logically straight forward, and as long as you don't mind calculating a lot of numbers you will get the answer so I have no doubt that you will get any of these wrong from what I have seen so far

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

can we go over the last two really quickly

OpenStudy (yuki):

but all of the problems here can be found a lot more easily using the fact P(A) = 1-P(notA)

OpenStudy (anonymous):

4C1 (1/2)^1(1/2)^3 + 4C0 (1/2)^0 (1/2)^4 ????

OpenStudy (yuki):

so d is 1 - P (all of them being prime) and e is P(0) + P(1)

OpenStudy (anonymous):

is that right? what I just put

OpenStudy (yuki):

nope

OpenStudy (anonymous):

what is it then

OpenStudy (yuki):

sorry, ask someone else since I need to go. :)

OpenStudy (anonymous):

????

OpenStudy (anonymous):

what is it then